OpenStudy (anonymous):

integrate sec^2(1-4x) please :) i'm doing horrible with integration

myininaya (myininaya):

do you know what the integral of sec^2(x) w.r.t x is?

OpenStudy (anonymous):

isn't it tan x?

myininaya (myininaya):

\[\int\limits_{}^{}\sec^2(x) dx=? \]

myininaya (myininaya):

yes tan(x)+c

myininaya (myininaya):

so just do a sub here let u=1-4x

OpenStudy (anonymous):

should this be (tanx/1-4x)+c?

OpenStudy (anonymous):

use the substitution u=1-4x

OpenStudy (anonymous):

ok 1 sec lol sec

OpenStudy (anonymous):

-tanx/4 + c?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

-tan(1-4x)/4 + c

OpenStudy (anonymous):

my bad is that right?

myininaya (myininaya):

looks good.

OpenStudy (anonymous):

yes ,right

OpenStudy (anonymous):

ah man integrating trig functions is tough :( :(

myininaya (myininaya):

Its not too bad. You just have to practice a bit.

OpenStudy (anonymous):

how about \[\int\limits_{}^{?}\frac{ x }{ (4x^2+3 }\]

myininaya (myininaya):

what is derivative of 4x^2+3 w.r.t x?

OpenStudy (anonymous):

cause i have nothing at all for that one. been working on it for like three days. and im too broke to pay for wolfram lol. so thank you very much guys i am super grateful

OpenStudy (anonymous):

no,it is easy ,you can use integration by parts,substitution,_________________

OpenStudy (anonymous):

4/3(x^3)+c

OpenStudy (anonymous):

wait

myininaya (myininaya):

look if you let u=4x^2+3 isn't du/dx=8x don't you have an x on top?

OpenStudy (anonymous):

4/3(x^3) + 3x + c

OpenStudy (anonymous):

yes there is an x on top

OpenStudy (anonymous):

but then i end up with x=sqrt(u+3)/4

OpenStudy (anonymous):

is it really that sloppy?

myininaya (myininaya):

\[\frac{1}{8} \int\limits_{}^{}\frac{8x}{4x^2+3} dx\]

myininaya (myininaya):

you should the derivative of the bottom on top so this has form \[\frac{1}{8} \int\limits_{}^{}\frac{du}{u}\]

OpenStudy (anonymous):

sorry!!! oh my. the (4x+3) is (4x+3)^6

OpenStudy (anonymous):

wait (4x^2+3)^6

OpenStudy (anonymous):

that x squared, or that 6th power, is defeating me. i could solve it if one of them were missing but alas, there they are

myininaya (myininaya):

then you have this form which still do -able \[\frac{1}{8} \int\limits_{}^{}\frac{du}{u^6}\]

OpenStudy (anonymous):

ok i will try that right now. one sec

OpenStudy (anonymous):

doesn't du=8xdx?

myininaya (myininaya):

yes

OpenStudy (anonymous):

so dx=1/8x du?

OpenStudy (anonymous):

what am i supposed to do with that extra x?

myininaya (myininaya):

there was an x on top already

myininaya (myininaya):

\[\int\limits_{}^{}\frac{x}{(4x^2+3)^6} dx \] u=4x^2+3 du=8x dx or du/8=xdx

myininaya (myininaya):

replace the x dx with du/8

OpenStudy (anonymous):

AHHHHH

OpenStudy (anonymous):

you just blew my mind!!!

myininaya (myininaya):

I just divided both sides of du=8x dx by 8

myininaya (myininaya):

to solve for x dx should be

myininaya (myininaya):

in terms of u

OpenStudy (anonymous):

i know it was so obvious in front of my face thank you so much

OpenStudy (anonymous):

this is for a test review, the test is next monday. the questions the professor gave us in the review are much tougher than in the book. there was no example like that in the book, with a fraction with exponent on x and the whole term in the denominator

OpenStudy (anonymous):

one sec, trying to solve it now

OpenStudy (anonymous):

so it should be 1/40(4x^2+3)^5?

OpenStudy (anonymous):

-1/40(4x^2+3)^5

OpenStudy (anonymous):

ah thank you guys so much

OpenStudy (anonymous):

how about for \[\int\limits_{}^{}\frac{ x^2 }{ x^3+5 }\], should i set u=x^3+5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k, that seemed wrong to me for some reason, but i will try it now

OpenStudy (anonymous):

ah im beginning to see it

OpenStudy (anonymous):

i end up with 1/3 int u^-1 du

myininaya (myininaya):

ok and that equals?

OpenStudy (anonymous):

1/3u^0du

OpenStudy (anonymous):

so would that be 1/3 (x^3+5)

myininaya (myininaya):

\[\frac{1}{3} \int\limits_{}^{}u^{-1} du \text{ or } \frac{1}{3}\int\limits_{}^{}\frac{1}{u} du \\ =\frac{1}{3}\ln|u|+C\]

OpenStudy (anonymous):

omg wow i'm so ashamed i didn't see that

OpenStudy (anonymous):

deriv and int of ln u are my faves :(

myininaya (myininaya):

yeah you can't use the power rule for when the n is -1

OpenStudy (anonymous):

yeah it ends up being u^0=1 right

myininaya (myininaya):

as in \[\int\limits_{}^{}u^{n} du=\frac{u^{n+1}}{n+1} +C , n \neq -1 \]

myininaya (myininaya):

well mainly because this fraction I just mentioned isn't defined at n=-1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

lol

myininaya (myininaya):

so when n=-1 we have \[\int\limits_{}^{}u^{-1} du =\ln|u|+c \\ \text{ and not } \frac{u^{0}}{0}+c\] because fractions aren't defined when their bottom is 0

OpenStudy (anonymous):

right, i forgot the n+1 in the denom

OpenStudy (anonymous):

what about for \[\int\limits_{}^{}\frac{ 1 }{ 1+(e^-2x) }\], can this be done with u=1+e^-2x?

OpenStudy (anonymous):

is that your ques\[\int\limits_{}^{}\frac{ 1 }{ 1+e^{-2x} }\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sorry, new to this site's format

OpenStudy (anonymous):

so can this be done with u sub?

myininaya (myininaya):

honestly this might require an algebraic manipulation for the algebraic substitution

myininaya (myininaya):

this might help (try getting rid of that neg exponent on bottom) multiply top and bottom by e^(2x)

OpenStudy (anonymous):

so i end up with e^2x/2?

myininaya (myininaya):

no

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1 \cdot (e^{2x})}{1 \cdot (e^{2x})+e^{-2x} \cdot (e^{2x})} dx\]

OpenStudy (anonymous):

you can multiply by e^(2x) :\[\frac{ 1 }{ 1+e^{-2x}} to get\frac{ e^{2x} }{ e^{2x}+1 }\],then use u=e^x

myininaya (myininaya):

i would actually use the sub u=e^(2x)+1

myininaya (myininaya):

but you can always do two subs too

OpenStudy (anonymous):

what if i did not algebraically manipulate it and used u=1+e^-2x? would it still work?

myininaya (myininaya):

if you differentiate that would you get the thing that is on top? (or some constant multiple times what you have on top)?

OpenStudy (anonymous):

no :( ok i see now

OpenStudy (anonymous):

is the answer ln|e^2x|?

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

du=2e^2xdx right?

OpenStudy (anonymous):

so e^2xdx=1/2du

OpenStudy (anonymous):

so it should be 1/2(ln|e^2x|) yes?

myininaya (myininaya):

\[\int\limits\limits_{}^{}\frac{1}{1+e^{-2x}} dx=\int\limits\limits_{}^{}\frac{e^{2x}}{e^{2x}} \cdot \frac{1}{1+e^{-2x}} dx =\int\limits\limits_{}^{}\frac{e^{2x}}{e^{2x}+1} dx \\ \text{ Let } u=e^{2x}+1 \\ \text{ so } du=2 e^{2x} dx \\ \text{ We see } e^{2x} \text{ on top so we decide to solve } du=2e^{2x} dx \\ \text{ for } e^{2x} dx \\ \text{ So this means we have } \frac{1}{2} du =e^{2x} dx \\ \int\limits\limits_{}^{}\frac{1}{2} \frac{1}{u} du =\frac{1}{2} \int\limits\limits_{}^{}\frac{1}{u} du =\frac{1}{2} \ln|u| +C \]

myininaya (myininaya):

And what was u again?

myininaya (myininaya):

You should be able to see that in my second line.

OpenStudy (anonymous):

e^2x+1

OpenStudy (anonymous):

oops. i forgot the +1

OpenStudy (anonymous):

SO MANY THINGS IN ADVANCED CALC ;_; MY POOR BRAIN